OpenStudy (anonymous):

Solve -6(5-2z)+5z=8z-3(16-3z)

OpenStudy (anonymous):

first multiply out using the distributive law

OpenStudy (anonymous):

-30+12z+5z=8z-42+9z

OpenStudy (anonymous):

then collect like terms

OpenStudy (anonymous):

\[-30+12z+5z=8z-42+9z\] \[-30+17x=17x-42\]

OpenStudy (mysesshou):

42 -->48

OpenStudy (anonymous):

then notice you have the same number of "z" on both sides

OpenStudy (anonymous):

so there is no way you can solve this, i.e. there is no solution

OpenStudy (anonymous):

and mysesshou is right, \[-3\times 16 = -48\] not what i wrote. still no solution

OpenStudy (anonymous):

so there is no answer to this its just another equation?

OpenStudy (mysesshou):

-6(5-2z)+5z=8z-3(16-3z) -30+12x+5z = 8z-48+9z -30+17z= 17z-48 Doesn't look like it. I get the same thing as satellite73

OpenStudy (anonymous):

i have been wrecking my brain for nothing lol

OpenStudy (mysesshou):

no matter what you substitute in for z, it is not equal \[\neq\]

OpenStudy (anonymous):

okay thanks guys